# Overcoming the Zero-Rate Hashing Bound with Holographic Quantum Error Correction

### v1

@@ III Results
In our biased-noise threshold profiling, we tested 43 distinct Pauli biases and have displayed them on the ternary diagram shown in Figure 2. Each point of the triangular plot shown represents a pure Pauli error channel. Subsequently, points on the interior of the plot can be read by following the grid lines provided to the boundaries; for example, the point to the lower left of the central depolarizing noise marker can be interpreted as having relative biases $\bar{r}=(r_{X},~{}r_{Y},~{}r_{Z})$ of $(0.25,~{}0.25,~{}0.5)$. At each bias point 10,000 Monte Carlo simulations were performed per threshold curve data point, per layer. We display commensurate threshold curve results for some select data points in Appendix B for the zero-rate HaPPY code.
@@
If we look beyond the pure 1-Pauli portion of the ternary plots, we note more subtle behavior of these codes: for example, it can be seen that modest threshold increases emerge for the Steane and $\llbracket 6,1,3\rrbracket$ codes as we move towards pure 2-Pauli noise, i.e., $XY$, $YZ$, and $XZ$ noise. These changes can be seen more clearly by examining the threshold behavior of the codes by tuning a bias parameter $\eta$. In previous work, this aim was accomplished by comparing various zero-rate codes against the hashing bound, a useful lower bound for quantifying code-capacity channel performance [26]. However, in contrast to the present work, all previous works have examined the pure 1-Pauli noise limit associated with the hashing bound, and have not delved into combinations of two types of Pauli errors at once, a condition known to occur in the square-lattice GKP code under a symmetric Gaussian random displacement noise channel [17, 18, 19, 20]. As such, we tune the $\eta$ parameter within the range $0\leq\eta\leq+\infty$, extending previous work past the regime indicated by $\eta\geq 0.5$ [26, 21, 25]. The hashing bound is formally defined as
@@
where $R$ represents an achievable rate $k/n$ for a random stabilizer code and $H(\bar{p})$ represents the Shannon entropy [10, 11]
$H(\bar{p})=-\sum_{i\in\{X,Y,Z\}}p_{i}\log{p_{i}}~{}.$ ) . (3)
@@
Table 1 displays several specific threshold points of interest in the ternary plots from Figure 3. Noise biases and their corresponding bias vectors are displayed, and figures in green and gold signify threshold data which either attain or exceed the hashing bound, or come with $2\%$ of achieving the bound, respectively. As is evidenced in each of the plots, the zero-rate HaPPY code clearly achieves the hashing bound for all pure 1-Pauli biases, as well as for finite biases in which $\eta>10$. Additionally, in the limits for which $\eta\to 0$, the HaPPY code thresholds closely trace out the hashing bound as pure 2-Pauli noise is approached. Generally, the behavior of the HaPPY code emulates in large part the behavior seen of the generalized toric code family [25]. The Steane code portrays a distinct trend: in the pure 1-Pauli noise limit, the code performs significantly worse than all other codes tested. Nonetheless, the zero-rate Steane code eclipses the hashing bound for depolarizing noise ($\eta=0.5$), while also approaching the bound for finite biases as we approach the pure 2-Pauli limit ($\eta=0$). It was shown in previous work [24] that the color code exhibits similar behavior in the pure 1-Pauli bias regime; as such, it is logical to infer that the holographic Steane code manifests similar behavior in these limits. Due to the fact that both codes are built from self-dual CSS codes [37], in the 1-Pauli limit, only one set of stabilizers gives information about the error, while the other stabilizers do not yield additional information for decoding. As such, the problem reduces to a classical linear code with a check matrix defined by only half of the symplectic check matrix in the quantum code. Notwithstanding, it is known that many code concatenation schemes based on graph states can slightly exceed the hashing bound; our results, up to the uncertainty margin given, show that the holographic Steane can at least match most of the codes discovered in [28].
@@
IV Conclusion
In this work we have shown strong evidence of extremely high threshold behavior for zero-rate holographic quantum codes in the presence of 1-Pauli and 2-Pauli noise channels, for pure and finitely-biased noise regimes. In doing so, we have tested the HaPPY; Steane; $\llbracket 6,1,3\rrbracket$; and SCF codes, finding that all of them admit threshold points for some Pauli noise bias which either attain or surpass the hashing bound. In addition to these codes, the “tailored” $\llbracket 7,1,3\rrbracket$ code from [68] was also tested; we found that the biased-noise ternary plot is identical to that of the HaPPY code. This result is in line with the conclusions from [25], which constructed generalized toric codes with performance surpassing the hashing bound in the pure 1-Pauli limit. In related work [55], an integer-optimization decoder was utilized for the zero-rate Evenbly code in order to pinpoint the depolarizing and pure 1-Pauli thresholds of $19.1\%$ and $50\%$, respectively. It would be interesting to see if a tensor-network decoder would allow for such complete biased-noise threshold explorations of Evenbly codes [69, 70, 71]. Our work demonstrates that holographic codes exhibit remarkable code-capacity properties under biased noise channels and can even overtake current state-of-the-art results [28]. It further indicates that they constitute a novel and competitive class of stabilizer codes that is robust against biased noise beyond the conventional options of topological codes [26, 23, 27, 24, 22, 21], naïve constructions of concatenated codes, or graph codes [28]. These results affirm that holographic quantum codes could have practical utility for quantum computing. Taking stock of our results and in combination with recent work showing that the bulk geometric transition in AdS/CFT implies a threshold for holographic codes in the continuum limit [61], it is natural to ask whether the discretized isometric map typical of holographic codes provides a general method for constructing codes which faithfully achieve various finite points of the hashing bound. We leave this direction for future work. In the interests of practical quantum computing, one application of holographic quantum codes may be found in magic-state distillation [72, 73]. Indeed, for the case of the HaPPY and holographic Steane codes, the transversal logical $\overline{SH}$ and $\bar{H}$ gates allow for current state-of-the-art methods to be leveraged, as was pointed out in [74]. Also, as the holographic map can be applied to XP-stabilizer codes [75, 44], we additionally expect that further examples of holographic quantum codes can be constructed and examined. As a final note, we have only tested the zero-rate versions of many holographic codes, but one may test constant-rate holographic codes as well, as shown in [34]; the construction and testing of such codes is the subject of active research.
@@ IV Conclusion
Acknowledgements. — We thank Aritra Sarkar, David Elkouss, and Jens Eisert for discussions related to concatenated codes, universal logical gate sets, as well as quantum channel capacity and the hashing bound. We also thank Mackenzie Shaw for pointing out that the square-lattice GKP code admits an induced $XZ$ 2-Pauli noise channel under a symmetric Gaussian random displacement noise model. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”.
@@
Author Contributions. — JF developed and implemented the tensor-network decoder during his master thesis, under the supervision of MS and SF. MS wrote the manuscript. AJ, CC, and SF provided project guidance.
Appendix A Seed Tensors for Other Holographic Codes
The seed tensors and their associated logical operators and their stabilizers are shown in Table 2. Here we have used the notation $XX$ to denote $X\otimes X$ for brevity. Additionally, the final entry for each stabilizer generator (with a bar above) denotes the action on the logical index.
Seed Tensor Stabilizers Logical Operators
Appendix B Threshold Results for Select Data Points
In Figure 5 we display the individual threshold curves obtained for the zero-rate HaPPY code for depolarizing noise, as well as for biased pure $X$, pure $Y$, and pure $Z$ noise. Around the fixed points in subfigures (b)-(d), the recovery rate promptly increases, which is due to several reasons: Firstly, the pure 1-Pauli noise capacity can be shown to be $50\%$, in agreement with known channel-capacity findings [10]. Secondly, as our tensor-network decoder works essentially as a maximum-likelihood decoder, after the $p=0.50$ mark, the decoder selects the most-probable error and returns a pure error vector which is used to correct. Therefore, unlike a minimum weight decoder, the tensor network decoder is still able to provide the right correction with high probability even when $p>0.50$. More details can be found in [30, 76].

### v2

@@ III.1 Noise Model and Setup
In our biased-noise threshold profiling, we tested 43 distinct Pauli biases and have displayed them on the ternary diagram shown in Figure 2. Each point of the triangular plot shown represents a pure Pauli error channel. Subsequently, points on the interior of the plot can be read by following the grid lines provided to the boundaries; for example, the point to the lower left of the central depolarizing noise marker can be interpreted as having relative biases $\bar{r}=(r_{X},~{}r_{Y},~{}r_{Z})$ of $(0.25,~{}0.25,~{}0.5)$. At each bias point 10,000 Monte Carlo simulations were performed per threshold curve data point, per layer. We display commensurate threshold curve results for some select data points in Figure 4 for the zero-rate HaPPY code.
III.2 Zero-Rate Holographic Code Results
@@
As a more detailed example, we plotted in Figure 4 the individual threshold curves obtained for the zero-rate HaPPY code for depolarizing noise, as well as for biased pure $X$, pure $Y$, and pure $Z$ noise. Around the fixed points in subfigures (b)-(d), the recovery rate promptly increases, which is due to several reasons: Firstly, the pure 1-Pauli noise capacity can be shown to be $50\%$. Secondly, as the tensor-network decoder is a maximum-likelihood decoder, after the $p=0.50$ mark, the decoder selects the most-probable error and returns a pure error (destabilizer) vector which is used to correct. Therefore, unlike a minimum weight decoder, the tensor network decoder is still able to provide the right correction with high probability even when $p>0.50$. More details can be found in [40, 72].
If we look beyond the pure 1-Pauli portion of the ternary plots, we note more subtle behavior of these codes: for example, it can be seen that modest threshold increases emerge for the Steane and $\llbracket 6,1,3\rrbracket$ codes as we move towards pure 2-Pauli noise, i.e., $XY$, $YZ$, and $XZ$ noise. These changes can be seen more clearly by examining the threshold behavior of the codes via the tuning of a bias parameter $\eta$. In previous work, this aim was accomplished by comparing various zero-rate codes against the hashing bound, a useful lower bound for quantifying code-capacity channel performance [25]. However, in contrast to the present work, all previous works have examined the pure 1-Pauli noise limit associated with the hashing bound, and have not delved into combinations of two types of Pauli errors at once, a condition known to occur in the square-lattice GKP code under a symmetric Gaussian random displacement noise channel [74, 75, 76, 77], as well as for error channels resulting from performing Pauli twirling on the amplitude damping noise channel [78]. As such, we tune the $\eta$ parameter within the range $0\leq\eta\leq+\infty$, extending previous work past the regime indicated by $\eta\geq 0.5$ [25, 20, 24].
The hashing bound is formally defined as
@@
where $R$ represents an achievable rate $k/n$ for a random stabilizer code and $H(\bar{p})$ represents the Shannon entropy [1]
$H(\bar{p})=-\sum_{i\in\{I,X,Y,Z\}}p_{i}\log{p_{i}}~{}.$ ) . (6)
@@
Table 2 displays several specific threshold points of interest in the ternary plots from Figure 3. Noise biases and their corresponding bias vectors are displayed, and figures in green and gold signify threshold data which either attain or exceed the hashing bound, or come within $2\%$ of achieving the bound, respectively.
As is evidenced in each of the plots, the zero-rate HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes clearly achieve the hashing bound for all pure 1-Pauli biases, as well as for finite biases in which $\eta>10$. Additionally, in the limits for which $\eta\to 0$, the HaPPY and tailored $\llbracket 7,1,3\rrbracket$ code thresholds closely trace out the hashing bound as pure 2-Pauli noise is approached. The behavior of the HaPPY code emulates in large part the behavior seen of the generalized toric code family [24]; we discuss this in more detail in Section IV.
@@
The Steane code portrays a distinct trend: in the pure 1-Pauli noise limit, the code performs significantly worse than all other codes tested. Nonetheless, the zero-rate Steane code eclipses the hashing bound for depolarizing noise ($\eta=0.5$), while also approaching the bound for finite biases as we approach the pure 2-Pauli limit ($\eta=0$). It was shown in previous work [23] that the color code exhibits similar behavior in the pure 1-Pauli bias regime; as such, it is logical to infer that the holographic Steane code manifests similar behavior in these limits. Due to the fact that both codes are built from self-dual CSS codes [42], in the 1-Pauli limit, only one set of stabilizers gives information about the error, while the other stabilizers do not yield additional information for decoding. As such, the problem reduces to a classical linear code with a check matrix defined by only half of the symplectic check matrix in the quantum code. Notwithstanding, it is known that many code concatenation schemes based on graph states can slightly exceed the hashing bound; our results, up to the uncertainty margin given, show that the holographic Steane can at least match most of the codes discovered in [27].
Figure 6: Threshold curves for the Clifford-deformed Steane code that we mention in Section IV. Here, we have applied Hadamard gates to physical qubits $2,3,5,$ and $6$. In (a) we document a slight threshold increase to $19.0\pm 0.33\%$. (b) and (d) show the thresholds under pure X and Z noise, respectively, under which an increase to approximately $50\%$ can be observed. Finally, (c) displays the results for pure Y noise, with a slight decrease to $10.58\pm 0.12\%$ apparent.
In [21, 20], it was shown that randomized Clifford deformations, when applied to the surface code, can improve threshold performance with respect to a code-capacity noise channel. In this way, a surface code may be tailored to achieve the 1-Pauli hashing bound; one wonders whether such randomized approaches could be leveraged in holographic codes as well. Indeed, we may regard the additional Hadamard gates in the Evenbly code [56, 79] as an instance of Clifford deformation, in order to uphold the strict isometry properties present, while utilizing less highly-entangled quantum states.
As a preliminary step in this direction, we have performed a random Clifford-deformation of the holographic Steane code using Hadamard gates, with the goal of improving the 1-Pauli threshold for the code. The seed tensors used in this modified holographic code are Clifford-equivalent to that of the original Steane code; the only difference lies in the application of Hadamard gates to physical qubits $2,3,5$, and $6$. The resulting stabilizer generators and logical operators for this Clifford-deformed Steane code are given in Table 1.
As is shown above, the stabilizers and logical operators do not follow the typical convention as expected for mixing Pauli operators in a cyclic manner, nor as in generalized toric code constructions [24, 38, 25, 20]. The thresholds, reported in Figure 6, are for: (a) depolarizing noise ($19.0\pm 0.33\%$); (b) pure X noise ($\sim 50\%$); (c) pure Y noise ($10.58\pm 0.12\%$); and (d) pure Z noise ($\sim 50\%$). We observe that even our naive Clifford deformations, when combined with the holographic concatenation method, can greatly improve the threshold performance for targeting noise models of interest. Indeed, our results show that the Clifford-deformed holographic Steane code achieves the hashing bound for pure X and Z noise, and supersedes the bound for depolarizing noise. Such modifications will be investigated and are the subject of active research.
@@
IV Discussion
In this work we have shown strong evidence of extremely high threshold behavior for zero-rate holographic codes in the presence of 1-Pauli and 2-Pauli noise channels, for pure and finitely-biased noise regimes. In doing so, we have tested the: HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes; Steane code; $\llbracket 6,1,3\rrbracket$ code; SCF code; and a Clifford-deformed Steane code. All of these codes admit thresholds for some Pauli noise bias which either attain or surpass the hashing bound, and one of these codes, the SCF, surpasses known records for 2-Pauli noise. Our work thus demonstrates that holographic codes exhibit remarkable code-capacity properties under biased noise channels and can even overtake current state-of-the-art results [27]. It moreover indicates that they constitute a novel and competitive class of stabilizer codes that is robust against biased noise, beyond the conventional options of topological codes [25, 22, 26, 23, 21, 20], naïve constructions of concatenated codes [80], or advanced code-concatenation techniques [27]. Our results heavily imply that further modification via Clifford deformations is possible, and additionally that holographic codes can in fact be tailored to Pauli noise channels of interest.
Our work leaves many open questions for the future. Firstly, we mentioned previously that the limit of 1-Pauli noise (that is, $\eta\mapsto\infty$) for both the HaPPY and tailored $\llbracket 7,1,3\rrbracket$ codes achieve the zero-rate hashing bound. The results that we have shown are in line with those of [24, 38], wherein such basic seed tensors were built up as in the structure of cyclic XZZX codes, and then mapped to generalized toric codes. Here, although the threshold values that we find are perfectly in line with the results of [24], our approach differs substantially in that we have uncovered a new class of subsystem stabilizer code for which several variants can achieve the zero-rate hashing bound for 1-Pauli noise, in line with several other known code constructions [27, 20, 25, 22, 23, 21].
In related work [56], an integer-optimization decoder was utilized for the zero-rate Evenbly code in order to pinpoint depolarizing and pure 1-Pauli noise thresholds of $19.1\%$ and $50\%$, respectively overcoming and achieving the hashing bound. It would be interesting to see if a tensor-network decoder would allow for a more complete biased-noise threshold exploration of Evenbly codes and their many derivatives [79, 81, 82]. However, in order to realize such an aim, the current tensor-network decoder utilized in our study would need to incorporate the extra Hadamard gates on edges, in order to extract the correct destabilizers from a syndrome. We leave such pursuits for subsequent work.
Additionally, it is known from [56, 79] that Evenbly codes exhibit gauge-dependent threshold behavior, owing to the specific isometric constraints used in the structure of the code. In principle, nothing restricts us from utilizing such gauge-fixing techniques in conjunction with maximum-rate constructions of the holographic codes studied here. However, the unique seed-tensor isometry properties of each holographic code may be utilized in order to give rise to similar gauge-dependent threshold “phases”, as the maximum-rate HaPPY code under gauge-fixing can exhibit very high threshold behavior [83] on a $\{5,5\}$ hyperbolic tiling.
Concerning the topic of finite-rate holographic code constructions, it is currently unknown whether there exist logical-qubit implantation schemes which maximize a code-capacity threshold for a given error model. It was shown in [37] that several naive logical-qubit implantation schemes can approach the hashing bound; furthermore, it has been conjectured in [56] that certain finite-rate conceptions of Evenbly codes may in fact achieve the code-capacity threshold for quantum erasure noise, as well as the finite-rate hashing bound for Pauli noise. Investigating these possibilities will be at the center of future research efforts.
Recently, generalizations of Abelian stabilizer codes have been investigated as alternatives for low-overhead quantum error correction, as such codes naturally allow for non-Clifford logical gates [50, 84, 85]. The quantum LEGO formalism itself is not restricted to only stabilizer circuits, and may be readily utilized for simulating aspects of either non-Abelian stabilizer or non-stabilizer codes. In this case, an enumerator-based method that combines tensor network and Monte Carlo sampling can be used to estimate logical error probabilities and thresholds of the code [41]. Alternatively, it is possible to calculate the coherent information instead [86, 87]. However, new techniques are likely needed to calculate the coherent information for such generalized holographic codes, since the number of physical qubits at the boundary grows exponentially with a commensurate number of layers [35, 54].
Comparing all of our results with those of concatenated codes, we note that the codes tested here fare better than typical tree-style concatenation [80, 88], particularly for the depolarizing noise channel. Nevertheless, it was shown in [27] that different types of repetition code concatenations can achieve threshold values comparable to ours for the depolarizing noise limit. Additionally, it was shown that various different concatenation schemes achieve or slightly surpass the hashing bound, depending on the details of the two repetition codes involved in the concatenation. Given this, one may then ask as to why our zero-rate holographic code constructions behave as they do. Considering the structural differences between holographic codes and more straightforward code-concatenation schemes, we would expect that in the 2-Pauli limit a conferrable advantage must be present by concatenation along the edges of a hyperbolic tessellation. Our forthcoming work will investigate this advantage more thoroughly.
As a final observation from the more theoretical perspective, recent work [65] has shown that the bulk geometric transition in AdS/CFT implies a threshold for holographic codes in the continuum limit. Taking this finding in tandem with our results, it is natural to ask whether the discretized isometric map typical of holographic codes provides a general method for achieving above-hashing-bound behavior, and additionally, whether or not AdS/CFT in the continuum limit implies the existence of a fault-tolerance threshold for holographic codes.
IV.1 Holographic Codes in Practice
In this work, we have investigated the threshold of holographic codes with respect to code-capacity noise channels. Although it has not been confirmed whether or not holographic codes admit a fault-tolerance threshold [65], it seems unreasonable to suspect otherwise, given the fact that the threshold data we have presented here is comparable to many state-of-the-art code families. A first step could be to test phenomenological noise models, i.e., those that incorporate measurement noise into the syndrome-extraction process. This goal could in principle be achieved by mapping the tensor network to a detector picture, as was recently performed in [89], thereby facilitating large-scale circuit-level noise simulations.
In the interim, one can consider techniques by which full fault-tolerant syndrome extraction and universal logic can be performed. [90, 91, 92] proved that generalized concatenated code schemes based on the quantum Hamming code can exhibit constant space overhead and quasi-polylogarithmic time overhead. These schemes, all based on code concatenation, involve the use of hard-decision layerwise decoding protocols [93]. As it has been proven that holographic quantum codes are specific instances of generalized concatenated codes for the hyperbolic plane [67], we would surmise that the aforementioned techniques can also be adapted to our setting. This idea is the subject of active research.
Many alternatives have been suggested as well to the layerwise decoding scheme. One method may involve floquetifying holographic codes [94]; this procedure would solve two problems at once, in that high-weight stabilizers and logical operators could be controlled to a reasonable weight, and fault-tolerance can be guaranteed with an appropriate measurement schedule. In addition to these ideas, fault-tolerant logic has also been shown to be gauge-dependent [56] and universal in some cases for holographic codes [67]. Recent work has also made progress on lowering the computational complexity of decoding itself using tensor-network schemes [95, 96]. These techniques, often based on the analysis of critical or approximate contraction paths, could allow for even faster decoding than is currently employed by polynomial-complexity tensor-network decoding [37, 36], and may become a prime factor in the construction practical quantum codes which supersede the hashing bound.
One may also ask what specific applications may be well-suited to the usage of a holographic code. An application of holographic codes may be found in magic-state distillation [97, 98]. Indeed, for the case of the HaPPY and holographic Steane codes, the transversal logical $\overline{SH}$ and $\bar{H}$ gates allow for current state-of-the-art methods to be leveraged, as was pointed out in [99]. For practical quantum error correction, it is known that the square-lattice GKP code admits an induced $XZ$ 2-Pauli noise channel under a symmetric Gaussian random displacement noise model [74, 75, 76, 77], and also that Pauli-twirling the amplitude damping channel results in biased 2-Pauli noise [78]. Such applications could be combined with our work in the future.
IV.2 Acknowledgements
@@
We thank Aritra Sarkar, David Elkouss, and Jens Eisert for discussions related to concatenated codes and quantum channel capacity. We also thank Mackenzie Shaw and Kathleen Chang for mentioning at the conference FTQC ’24 in Benasque several realistic examples of 2-Pauli noise arising in practical quantum computing. MS and SF acknowledge financial support from the Intel corporation. AJ is supported by the Einstein Research Unit “Perspectives of a quantum digital transformation”.
IV.3 Author Contributions
@@
JF developed and implemented the tensor-network decoder during his master thesis, under the supervision of MS and SF. MS, CC, and AJ wrote the manuscript. AJ, CC, and SF provided project guidance.